Circular No. 6080

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1915+105 S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on

behalf of the Granat WATCH team, communicate: “We are currently observing a strong outburst from GRS 1915+105. Preliminary average flux estimates for Sept. 14.5, 17, 18, and 19 UT are 780 ± 40, 460 ± 50, 560 ± 30, and 830 ± 40 mCrab, respectively, in the energy band 8–20 keV. The previous WATCH observation of GRS 1915+105 in its high state was in March (IAUC 5959), when the flux reached 400 mCrab and a superluminal radio source appeared, according to Mirabel and Rodriguez (1994, Nature 371, 46).”

N. Alexandrovich, K. Borozdin, and R. Sunyaev, on behalf of the MIR-KVANT-TTM team, report: “Strong x-ray flux from GRS 1915+105 was registered by the TTM coded-mask instrument onboard the MIR space station during observations on Sept. 8 (500 mCrab at 5 keV), 10 (460 mCrab), and 11 (840 mCrab). The spectrum of the source is very soft (black-body fit gives the radiation temperature as 1.25 ± 0.05 keV with absorption column N_{H} = 2 × 10^{22} cm^{−2}). The observation of Sept. 11 showed the appearance of the hard tail with flux 140 ± 30 mCrab in the spectral band 17–25 keV, in addition to the soft component.”

X-RAY NOVA IN SCORPIUS R. Gonzalez, University of California at Berkeley; D. Martins, Univer-

sity of Alaska; and C. Bailyn and J. Orosz, Yale University, write: “The recent secondary outburst of GRO 1655–40 (IAUC 6075, 6077, 6079) does not appear to have resulted in substantial changes in the V -band magnitude of the optical counterpart, as shown by the following CCD observations obtained with the Cerro Tololo 0.9-m reflector: Sept. 9.046 UT, V = 15.09; 10.041, 14.97; 11.032, 15.05; 15.076, V = 15.19, V − I = +1.93. Further monitoring of the decline of this object is strongly encouraged.”

SATURN Further to their comments on IAUC 6079, A. Sanchez-Lavega et al.

note that the spot reported on IAUC 6059 at ∼ −60^{o} has expanded toward the east while becoming less reflective; on Aug. 28.1, it had increased its length to ∼ 50^{o}, retaining a bright nucleus of ∼ 5^{o} that drifts −3^{o}.5/day relative to System III.

1994 September 21 Daniel W. E. Green